if t is in quadrant II and csc t=2, find the value of cot t.
Found 3 solutions by Alan3354, solver91311, jrfrunner:
Answer by Alan3354(69443)

if t is in quadrant II and csc t=2, find the value of cot t.
-----------
If csc(t) = 2, sin(t) = 1/2.
arcsin(1/2) = 150 degs in Quad II.
cot(150) = - sqrt(3)

Answer by solver91311(24713)



in Quadrant II

which is to say:



In Quadrant II, and



Hence

and

Furthermore:



So



But for Quadrant II so we need

Next:



Hence:




John

My calculator said it, I believe it, that settles it


Answer by jrfrunner(365)

Quadrant II is where x=cost is negative and y=sint is positive
---
csc(t) = 2; 1/sin(t)=2; sin(t)=1/2 therefore t=30 (actually 180-30=150)
--
Cot(t)=cos(t)/sin(t)
cot(150)=cos(150)/sin(150) = -cos(30)/sin(30)=(-sqrt(3)/2)/(1/2)=-sqrt(3)